Green beatnotes recovered.
It was just a matter of aligning the arm greens and PSL greens on the PSL table. I suppose something knocked the PSL alignment out of whack... I was also able to simultaneously see the green beatnote and IR beatnote respond to Yend laser temperature.
Locked arms on POX/POY, checked RMS of ALS-BEAT[X/Y]_FINE_PHASE_OUT_HZ channels.
These seem fine. Locked CARM and DARM on ALS, found IR resonances.
ALS is back in business